OpenStudy (anonymous):

MEDAL AND FAN Rewrite f(x) = –2(x − 3)^2 + 2 from vertex form to standard form

OpenStudy (jhannybean):

First define what vertex and standard form is: vertex form : \(\sf y= a(x-h)^2+k\) Standard form: \(\sf y=ax^2+bx+c\)

OpenStudy (jhannybean):

k and c are both constants.

OpenStudy (anonymous):

-2x^2+12x+20

OpenStudy (jhannybean):

\[(x-3)^2 = x^2-6x+9 \]\[-2(x^2-6x+9) = -2x^2+12x+18+2\]\[-2x^2+12x+20 ~\checkmark\] Correct.

OpenStudy (anonymous):

are u postive